ADNAN MUHAMMED

Grade 12,

Ques.1> A capacitor of capacitance 10microfarad is charged to a potential of 1V. It is connected to an inductor of inductance 10 raise to power -3 H. The maximum current that will flow in the circuit has the value - a> 1 micro ampere b> 1 milli ampere c> under root 1000 milli ampere d> 100 milli Ampere Ques 2> In a region of uniform magnetic induction B = 10 raise to the power -2 T, a circulaer coil of radius 30cm and resistance pie square ohm is rotatted about an axis which is diameter of the coil. If the coil rotates at 200rpm, the amplitude of the alternating current induced in the coil is a> 4 pie square milli ampere b> 30 milli ampere c> 6 milli ampere d> 200 milli ampere Ques.3> A circular coil of raidus 0.1 m has 80 turns of wire. If the magnetuc field thorugh the coil increases from 0 to 2 tesla in 0.4 sec and the coil is connected to a 11 ohm resistor, what is the current (in A) through the resisitor during the 0.4 sec? a> (8/7) A b> (7/8) A c> 8 A d> 7 A Please help me with complete solutions need them URGENTLY.. Thanks in advance.. will rate u for help!! Rohit


Ques.1>  A capacitor of capacitance 10microfarad is charged to a potential of 1V. It is connected to an inductor of inductance 10 raise to power -3 H. The maximum current that will flow in the circuit has the value -
 
a> 1 micro ampere
b> 1 milli ampere
c> under root 1000 milli ampere
d> 100 milli Ampere
 
Ques 2> In a region of uniform magnetic induction B = 10 raise to the power -2 T, a circulaer coil of radius 30cm and resistance pie square ohm is rotatted about an axis which is diameter of the coil. If the coil rotates at 200rpm, the amplitude of the alternating current induced in the coil is
 
a> 4 pie square milli ampere
b> 30 milli ampere
c> 6 milli ampere
d> 200 milli ampere
 
Ques.3> A circular coil of raidus 0.1 m has 80 turns of wire. If the magnetuc field thorugh the coil increases from 0 to 2 tesla in 0.4 sec and the coil is connected to a 11 ohm resistor, what is the current (in A) through the resisitor during the 0.4 sec?
 
a> (8/7) A
b> (7/8) A
c> 8 A
d> 7 A
 
Please help me with complete solutions need them URGENTLY..
 
Thanks in advance.. will rate u for help!!
Rohit


Grade:10

1 Answers

ROSHAN MUJEEB
3 years ago
Time period for oscillating magnet is
T=2πMBI​.........(1)
where I is the moment of inertia,

M is the magnetic moment
M=x×l
x=polestrength
When the magnet is broken in two equal parts, the new magnetic dipole moment is
M′=2x×l​=2M​
I′=2ML12​
=122M​×(2L​)2​
=12×8ML2​
=8I​
New time period is
T′=2πM′BI′​
TT′​=M′I′​
Putting all the values we get
TT′​=41​​=21​